# The Bose-Hubbard polaron from weak to strong coupling

### v2

@@ Abstract
We investigate the zero-temperature properties of a mobile impurity immersed in a bath of bosonic particles confined to a square lattice. We analyze the regimes of attractive and repulsive coupling between the impurity and the bath particles for different strengths of boson-boson interactions in the bath, using exact large-scale quantum Monte-Carlo simulations in the grand canonical ensemble. For weak coupling, the polaron mass ratio is found to decrease around the Mott insulator (MI) to superfluid (SF) transition of the bath, as predicted by recent theory [1], confirming the possible use of the impurity as a probe for the transition. For strong coupling in the MI regime, instead, the impurity is found to modify the bath density by binding to an extra bath particle or a hole, depending on the sign of the polaron-bath interactions. While the binding prevent the aforementioned use of the polaron mass ratio as an MI-SF transition probe, we show that it can be used instead as a probe of the binding itself. Our exact numerical results provide a benchmark for comparing lattice Bose polaron theories and are relevant for experiments with cold atoms trapped in optical lattices, where the presence of a confining harmonic potential can be modeled by a slowly varying local chemical potential.
@@
A mobile impurity interacting with a bath can give rise to a quasiparticle known as a polaron [2, 3], a paradigmatic open quantum system relevant to semiconductors [4], superfluid helium [5], and nuclear matter [6]. Recent advances in ultracold atom experiments, including quantum gas microscopy [7, 8, 9, 10], have enabled detailed studies of impurities in Bose-Einstein condensates [11], both in the continuum and in optical lattices described by the Bose-Hubbard (BH) model [12, 13, 14, 15, 9, 16, 17, 18, 19, 20, 21]. This has opened up the study of polaron physics in the strongly interacting regime and in the vicinity of quantum phase transitions. Lattice Bose polarons have been recently explored near the Mott insulator–superfluid (MI–SF) transition using beyond-mean-field Quantum-Gutzwiller (QGW) methods [1], and in combination with diagrammatic and variational quantum Monte Carlo techniques [22] at fixed particle density and in the canonical ensemble. These works demonstrate that for sufficiently weak impurity-bath interactions, polaron spectral properties such as the energy shift and effective mass can serve as sensitive probes of the MI–SF transition and its universality class [1]. At strong coupling, novel bound states involving the impurity and bath particles-holes excitations have been shown to emerge for fixed atoms numbers, beyond the polaron picture [22]. A key challenge remains in providing exact results for polaron dynamics across all regimes of interaction, in particular for strong impurity-bath and intra-bath interactions and in the grand canonical ensemble. The latter is particularly interesting for neutral atom experiments, where the presence of external harmonic confinement for the atoms effectively realizes a situation where local atomic densities vary across the lattice, providing effective particle reservoirs.
@@
In this work, we use exact large scale quantum Monte Carlo (QMC) simulations based on the worm algorithm [23] to study a mobile impurity immersed in a Bose-Hubbard bath in two dimensions. Our multi-species worm algorithm (see Supplemental Material) can describe all regimes of impurity-bath interactions — from weak to strong, attractive to repulsive — including possible impurity-atom and impurity-hole binding. In order to characterize different regimes of impurity-bath interactions and possible new binding mechanisms, we focus on spectral properties such as the polaron effective mass and the bath-induced shift as well as impurity-bath correlation functions. For weak impurity-bath interactions, the polaron mass ratio is found to decrease around the MI-SF transition and the polaron can be used as a probe of the transition, as predicted by [1]. We quantify for what strength of interactions this picture ceases to describe the system dynamics and demonstrate that for strong interactions the impurity locally modifies the bath density resulting in a binding to an extra bath particle or a hole, depending on the sign and strength of the impurity-bath interactions (see Fig. 1(a)). These binding mechanisms, which tend to suppress the polaron mass ratio and differ from the case of constant density [22], produce distinct spectroscopic signatures. We discuss the connection of these results in the grand-canonical ensemble with neutral atom experiments in the presence of an external harmonic potential.
@@
We begin by analyzing the polaron at $\mu=(\sqrt{2}-1)U$, which corresponds to the mean-field tip of the first Mott lobe and allows for direct comparison with recent theoretical studies [1, 22]. For reference, the MI–SF transition in the 2D Bose–Hubbard model at unit filling along the commensurate density line has been determined via QMC to occur at $2dt/U=0.2389(6)$ [29]. In contrast to variational QMC results obtained in the canonical ensemble [22], our numerically exact simulations in the grand-canonical ensemble (see Supplemental Material) show that a single impurity does not shift the transition point or affect the Mott phase properties—such as the excitation gap or compressibility—in the thermodynamic limit.
@@
We characterize the quasiparticle properties of the polaron by analyzing its energy dispersion at small momenta, $E_{p}(\boldsymbol{k})=E_{0}+\frac{\boldsymbol{k}^{2}}{2M^{*}}+\mathcal{O}(\boldsymbol{k}^{4})$. Here, $E_{0}=E_{p}(\boldsymbol{k}=0)$ denotes the interaction-induced energy shift, while $M^{*}$ is the polaron’s effective mass, reflecting its mobility within the bath. To estimate $M^{*}$, we compute the impurity’s diffusion coefficient $D$ from its mean-squared displacement between imaginary times $\tau=0$ and $\tau=\beta/2$, and use the the relation $M^{*}=1/(2D)$ [30, 31]. Specifically,
@@
Impurity spectral properties and bound states— Figure 2 shows QMC data for the bath-induced energy shift $E_{0}$ (panels (a-c)), the onsite impurity-bath correlation $C_{IB}(0)$ (insets), and the polaron mass ratio $M/M^{*}$ (panels (d-f)), with $M=1/(2t)$. Results are shown as a function of $U_{IB}/U$ (both attractive and repulsive) for three bath regimes: deep in the MI phase (panels (a), (d)), within the MI phase but close to the MI-SF transition (panels (b), (e)), and in the SF phase (panels (c), (f)). In particular, panel (a) shows that, for a bath in the deep MI phase, the bath-induced shift $E_{0}$ takes the values $E_{0}\simeq E_{+1}=2U_{IB}+U-\mu$ (dashed dotted line), $U_{IB}$ (dashed line) and $\mu$ (dotted line) for increasing $U_{IB}/U$ in the regions $U_{IB}<\mu-U$, $\mu-U<U_{IB}<\mu$, and $U_{IB}>\mu$, respectively. These regions correspond to the configurations shown in Fig. 1(a): an impurity bound to an extra boson ($\text{MI}_{+1}$, left), a free polaron (MI, center), and an impurity bound to a hole ($\text{MI}_{-1}$, right). In particular, for small values of $U_{IB}/U$ in the MI phase, $E_{0}\simeq U_{IB}$, consistent with a mean-field picture of the impurity interacting with a homogeneous bath at $\langle n_{B}\rangle=1$. For strong repulsion, $E_{0}\simeq\mu$ reflects the energy cost of removing a bath particle, while for strong attraction, $E_{0}\simeq 2U_{IB}+U-\mu$ accounts for the addition of a boson and its interaction with the impurity and the existing bath particle. The picture above is corroborated by $C_{IB}(0)$, which varies from +1 to 0 to -1 across $\text{MI}_{+1}$, MI and $\text{MI}_{-1}$ regions, respectively. Notably, the extra hole or particle on the impurity site reflects a true change in the bath particle number, rather than a particle–hole pair excitation (see Supplemental Material). This contrasts with Ref. [22], where the impurity cannot alter the bath occupation.
@@
Panel (b) shows qualitatively similar features to (a), except for small deviations from the values $E_{+1}$, $U_{IB}$ and $\mu$ near the transition points between the different regions $\text{MI}_{+1}$, MI, and $\text{MI}_{-1}$. The transitions are instead completely smoothed out for the case of a weakly interacting bath in the SF regime (panel (c)), where $E_{0}/U$ increases smoothly and monotonically from attractive to repulsive $U_{IB}/U$. The values of $E_{0}/U$ are qualitatively accounted for by a continuous change in bath density on the site of the impurity and in the polaron cloud in this regime (see Supplemental Material).
@@
The three regions are also clearly identified from the mass ratio $M/M^{*}$ in panel (d-f). In particular, for a bath prepared well into the MI phase (panel (d)) and for $0.5\lesssim U_{IB}/U\lesssim 0.5$, $M/M^{*}$ is close to unity, consistent with the results of [1] for small $U_{IB}$. In the $\text{MI}_{\pm 1}$ regions, instead $M/M^{*}\simeq 0$, which is consistent with our picture above of an impurity bound to a particle and a hole, respectively. Indeed, on a 2D lattice, a free dimer consisting of two similar bosons bound by $U_{IB}$ is expected to have an effective hopping energy $2t^{2}/|U_{IB}|$ for $|U_{IB}|\gg t$ [32, 33], resulting in an expected $M/M^{*}\simeq 10^{-2}$, which is consistent with our QMC results. For panel (e), the transitions are again smoothed out, and the mass ratio is higher in the $\text{MI}_{\pm 1}$ regions, as expected from the higher value of $2dt/U$. Panel (f) shows a continuous evolution of $M/M^{*}$ from strong attractive to strong repulsive interactions for a weakly interacting bath in the SF phase. The characterization of impurity spectral properties across all interaction regimes and in particular the demonstration of dimer states in the $\text{MI}_{\pm 1}$ regimes in the grand canonical ensemble are main results of this work.
@@
MI-SF transition— Figure 3 shows the polaron mass ratio $M/M^{*}$ versus $2dt/U$ for various values of $U_{IB}/U$. In the weak coupling limit — darker circles and squares in the attractive (panel (a)) and repulsive (panel (b)) branches — the ratio stays close to unity, consistent with the picture of a mean-field-type free polaron deep inside the MI phase. As $2dt/U$ approaches the MI-SF boundary (vertical grey dashed line), $M/M^{*}$ exhibits a pronounced dip (see Inset), whose depth increases with $|U_{IB}/U|$. Beyond the transition, the mass ratio rises again in the SF phase. This non-monotonic “dip-rebound” behavior is consistent with quantum Gutzwiller (QGW) predictions [1], and serves as a clear spectral signature of the MI-SF transition in the weak coupling regime.
@@
As $|U_{IB}/U|$ increases beyond this regime, the polaron evolves from a mobile quasiparticle into a bound dimer in the MI phase. The mass ratio $M/M^{*}$ then displays two distinct behaviors depending on $2dt/U$. At low hopping, the impurity binds strongly to a particle or hole excitation in the bath, forming a heavy dimer-like state. This is reflected in the strongly suppressed mass ratio observed for $|U_{IB}/U|\gtrsim 0.6$ at low $2dt/U$, signaling the $\text{MI}_{\pm 1}$ region. As $2dt/U$ increases, the bound state destabilizes, and the system crosses over into the conventional MI phase, where the impurity regains mobility. This crossover, marked by curved dashed black line in the figure, occurs at progressively larger $2dt/U$ as the $|U_{IB}/U|$ increases, indicating a gradual strengthening of the bound state. This interpretation is further supported by the accompanying changes in bath density and same-site correlation $C_{IB}(0)$ (see Supplemental Material). These results show that the mass ratio ceases to serve as a reliable probe of the MI–SF transition in the strong-coupling regime. However, it remains a robust indicator of impurity–bath bound-state formation.
@@
In conclusion, we demonstrated through exact numerical analysis that increasing the impurity-bath coupling strength in a lattice Bose polaron—defined by a single impurity boson interacting with a Bose-Hubbard bath in a grand canonical ensemble to the emergence of an MI-like phase with incommensurate density. This occurs for both attractive and repulsive interactions, as the impurity binds an additional bath particle or hole. Moreover, we demonstrated that the properties of the polaron, such as its effective mass, exhibit distinct behaviors near the MI-SF phase transition, depending on whether the impurity-bath coupling is weak or strong. We showed that the polaron mass ratio can either be used as a probe for the bath’s MI-SF phase transition for weak coupling, or as a probe of the $\text{MI}_{\pm 1}$ bound state for strong coupling. We argued that these results are relevant for the experimental study of the Bose lattice polaron in an harmonic trap, where the added spatial complexity make those effects possible even for a fixed number of bosons and weak or strong impurity-bath interaction. This work opens several research directions, including the study of bipolaron formation and many-polaron problems in the strong interacting regime.

### v3

@@ Abstract
We investigate the zero-temperature properties of a mobile impurity immersed in a bath of bosonic particles confined to a square lattice. We analyze the regimes of attractive and repulsive coupling between the impurity and the bath particles for different strengths of boson-boson interactions in the bath, using exact large-scale quantum Monte-Carlo simulations in the grand canonical ensemble. For weak coupling, the polaron mass ratio is found to decrease around the Mott insulator (MI) to superfluid (SF) transition of the bath, as predicted by recent theory, confirming the possible use of the impurity as a probe for the transition. For strong coupling in the MI regime, instead, the impurity is found to modify the bath density by binding to an extra bath particle or a hole, depending on the sign of the polaron-bath interactions. While the binding prevent the aforementioned use of the polaron mass ratio as an MI-SF transition probe, we show that it can be used instead as a probe of the binding itself. Our exact numerical results provide a benchmark for comparing lattice Bose polaron theories and are relevant for experiments with cold atoms trapped in optical lattices, where the presence of a confining harmonic potential can be modeled by a slowly varying local chemical potential.
@@
A mobile impurity interacting with a bath can give rise to a quasiparticle known as a polaron [1, 2], a paradigmatic open quantum system relevant to semiconductors [3], superfluid helium [4], and nuclear matter [5]. Recent advances in ultracold atom experiments, including quantum gas microscopy [6, 7, 8, 9], have enabled detailed studies of impurities in Bose-Einstein condensates [10], both in the continuum and in optical lattices described by the Bose-Hubbard (BH) model [11, 12, 13, 14, 8, 15, 16, 17, 18, 19, 20]. This has opened up the study of polaron physics in the strongly interacting regime and in the vicinity of quantum phase transitions. Lattice Bose polarons have been recently explored near the Mott insulator–superfluid (MI–SF) transition using beyond-mean-field Quantum-Gutzwiller (QGW) methods [21], and in combination with diagrammatic and variational quantum Monte Carlo techniques [22] at fixed particle density and in the canonical ensemble. These works demonstrate that for sufficiently weak impurity-bath interactions, polaron properties such as the energy shift and effective mass can serve as sensitive probes of the MI–SF transition and its universality class [21]. At strong coupling, novel bound states involving the impurity and bath particles-holes excitations have been shown to emerge for fixed atoms numbers, beyond the polaron picture [22]. A key challenge remains in providing exact results for polaron properties across all regimes of interaction, in particular for strong impurity-bath and intra-bath interactions and in the grand canonical ensemble. The latter is particularly interesting for neutral atom experiments, where the presence of external harmonic confinement for the atoms effectively realizes a situation where local atomic densities vary across the lattice, providing effective particle reservoirs.
@@
In this work, we use exact large scale quantum Monte Carlo (QMC) simulations based on the worm algorithm [23] to study a mobile impurity immersed in a Bose-Hubbard bath in two dimensions. Our multi-species worm algorithm (for details see Supplemental Material [24]) can describe all regimes of impurity-bath interactions — from weak to strong, attractive to repulsive — including possible impurity-atom and impurity-hole binding. In order to characterize different regimes of impurity-bath interactions and possible new binding mechanisms, we focus on properties such as the polaron effective mass and the bath-induced shift as well as impurity-bath correlation functions. For weak impurity-bath interactions, the polaron mass ratio is found to decrease around the MI-SF transition and the polaron can be used as a probe of the transition, as predicted by [21]. We quantify for what strength of interactions this picture ceases to describe the system dynamic and demonstrate that for strong interactions the impurity locally modifies the bath density resulting in a binding to an extra bath particle or a hole, depending on the sign and strength of the impurity-bath interactions (see Fig. 1(a)). These binding mechanisms, which tend to suppress the polaron mass ratio and differ from the case of constant density [22], produce distinct signatures. We discuss the connection of these results in the grand-canonical ensemble with neutral atom experiments in the presence of an external harmonic potential.
@@
We begin by analyzing the polaron at $\mu=(\sqrt{2}-1)U$, which corresponds to the mean-field tip of the first Mott lobe and allows for direct comparison with recent theoretical studies [21, 22]. For reference, the MI–SF transition in the 2D Bose–Hubbard model at unit filling along the commensurate density line has been determined via QMC to occur at $2dt/U=0.2389(6)$ [30]. In contrast to variational QMC results obtained in the canonical ensemble [22], our numerically exact simulations in the grand-canonical ensemble [24] show that a single impurity does not shift the transition point or affect the Mott phase properties—such as the excitation gap or compressibility—in the thermodynamic limit.
@@
We characterize the quasiparticle properties of the polaron by analyzing its energy dispersion at small momenta, $E_{p}(\boldsymbol{k})=E_{0}+\frac{\boldsymbol{k}^{2}}{2M^{*}}+\mathcal{O}(\boldsymbol{k}^{4})$. Here, $E_{0}=E_{p}(\boldsymbol{k}=0)$ denotes the interaction-induced energy shift, while $M^{*}$ is the polaron’s effective mass, reflecting its mobility within the bath. To estimate $M^{*}$, we compute the impurity’s diffusion coefficient $D$ from its mean-squared displacement between imaginary times $\tau=0$ and $\tau=\beta/2$, and use the relation $M^{*}=1/(2D)$ [31, 32]. Specifically,
@@
Polaron properties and bound states— Figure 2 shows QMC data for the bath-induced energy shift $E_{0}$ (panels (a-c)), the onsite impurity-bath correlation $C_{IB}(0)$ (insets), and the polaron mass ratio $M/M^{*}$ (panels (d-f)), with $M=1/(2t)$. Results are shown as a function of $U_{IB}/U$ (both attractive and repulsive) for three bath regimes: deep in the MI phase (panels (a), (d)), within the MI phase but close to the MI-SF transition (panels (b), (e)), and in the SF phase (panels (c), (f)). In particular, panel (a) shows that, for a bath in the deep MI phase, the bath-induced shift $E_{0}$ takes the values $E_{0}\simeq E_{+1}=2U_{IB}+U-\mu$ (dashed dotted line), $U_{IB}$ (dashed line) and $\mu$ (dotted line) for increasing $U_{IB}/U$ in the regions $U_{IB}<\mu-U$, $\mu-U<U_{IB}<\mu$, and $U_{IB}>\mu$, respectively. These regions correspond to the configurations shown in Fig. 1(a): an impurity bound to an extra boson ($\text{MI}_{+1}$, left), a free polaron (MI, center), and an impurity bound to a hole ($\text{MI}_{-1}$, right). In particular, for small values of $U_{IB}/U$ in the MI phase, $E_{0}\simeq U_{IB}$, consistent with a mean-field picture of the impurity interacting with a homogeneous bath at $\langle n_{B}\rangle=1$. For strong repulsion, $E_{0}\simeq\mu$ reflects the energy cost of removing a bath particle, while for strong attraction, $E_{0}\simeq 2U_{IB}+U-\mu$ accounts for the addition of a boson and its interaction with the impurity and the existing bath particle. The picture above is corroborated by $C_{IB}(0)$, which varies from +1 to 0 to -1 across $\text{MI}_{+1}$, MI and $\text{MI}_{-1}$ regions, respectively. Notably, the extra hole or particle on the impurity site reflects a true change in the bath particle number, rather than a particle–hole pair excitation [24]. This contrasts with Ref. [22], where the impurity cannot alter the bath occupation.
@@
Panel (b) shows qualitatively similar features to (a), except for small deviations from the values $E_{+1}$, $U_{IB}$ and $\mu$ near the transition points between the different regions $\text{MI}_{+1}$, MI, and $\text{MI}_{-1}$. The transitions are instead completely smoothed out for the case of a weakly interacting bath in the SF regime (panel (c)), where $E_{0}/U$ increases smoothly and monotonically from attractive to repulsive $U_{IB}/U$. The values of $E_{0}/U$ are qualitatively accounted for by a continuous change in bath density on the site of the impurity and in the polaron cloud in this regime [24].
@@
The three regions are also clearly identified from the mass ratio $M/M^{*}$ in panel (d-f). In particular, for a bath prepared well into the MI phase (panel (d)) and for $0.5\lesssim U_{IB}/U\lesssim 0.5$, $M/M^{*}$ is close to unity, consistent with the results of [21] for small $U_{IB}$. In the $\text{MI}_{\pm 1}$ regions, instead $M/M^{*}\simeq 0$, which is consistent with our picture above of an impurity bound to a particle and a hole, respectively. Indeed, on a 2D lattice, a free dimer consisting of two similar bosons bound by $U_{IB}$ is expected to have an effective hopping energy $2t^{2}/|U_{IB}|$ for $|U_{IB}|\gg t$ [33, 34], resulting in an expected $M/M^{*}\simeq 10^{-2}$, which is consistent with our QMC results. For panel (e), the transitions are again smoothed out, and the mass ratio is higher in the $\text{MI}_{\pm 1}$ regions, as expected from the higher value of $2dt/U$. Panel (f) shows a continuous evolution of $M/M^{*}$ from strong attractive to strong repulsive interactions for a weakly interacting bath in the SF phase. The characterization of the polaron properties across all interaction regimes and in particular the demonstration of dimer states in the $\text{MI}_{\pm 1}$ regimes in the grand canonical ensemble are main results of this work.
@@
MI-SF transition— Figure 3 shows the polaron mass ratio $M/M^{*}$ versus $2dt/U$ for various values of $U_{IB}/U$. In the weak coupling limit — darker circles and squares in the attractive (panel (a)) and repulsive (panel (b)) branches — the ratio stays close to unity, consistent with the picture of a mean-field-type free polaron deep inside the MI phase. As $2dt/U$ approaches the MI-SF boundary (vertical grey dashed line), $M/M^{*}$ exhibits a pronounced dip (see Inset), whose depth increases with $|U_{IB}/U|$. Beyond the transition, the mass ratio rises again in the SF phase. This non-monotonic “dip-rebound” behavior is consistent with quantum Gutzwiller (QGW) predictions [21], and serves as a clear signature of the MI-SF transition in the weak coupling regime.
@@
As $|U_{IB}/U|$ increases beyond this regime, the polaron evolves from a mobile quasiparticle into a bound dimer in the MI phase. The mass ratio $M/M^{*}$ then displays two distinct behaviors depending on $2dt/U$. At low hopping, the impurity binds strongly to a particle or hole excitation in the bath, forming a heavy dimer-like state. This is reflected in the strongly suppressed mass ratio observed for $|U_{IB}/U|\gtrsim 0.6$ at low $2dt/U$, signaling the $\text{MI}_{\pm 1}$ region. As $2dt/U$ increases, the bound state destabilizes, and the system crosses over into the conventional MI phase, where the impurity regains mobility. This crossover, marked by curved dashed black line in the figure, occurs at progressively larger $2dt/U$ as the $|U_{IB}/U|$ increases, indicating a gradual strengthening of the bound state. This interpretation is further supported by the accompanying changes in bath density and same-site correlation $C_{IB}(0)$ [24]. These results show that the mass ratio ceases to serve as a reliable probe of the MI–SF transition in the strong-coupling regime. However, it remains a robust indicator of impurity–bath bound-state formation.
@@
In conclusion, we demonstrated through exact numerical analysis that increasing the impurity-bath coupling strength in a lattice Bose polaron—defined as a single impurity boson interacting with a Bose-Hubbard bath in a grand canonical ensemble—can give rise to the emergence of an MI-like phase with incommensurate density. This occurs for both attractive and repulsive interactions, as the impurity binds an additional bath particle or hole. Moreover, we demonstrated that the properties of the polaron, such as its effective mass, exhibit distinct behaviors near the MI-SF phase transition, depending on whether the impurity-bath coupling is weak or strong. We showed that the polaron mass ratio can either be used as a probe for the bath’s MI-SF phase transition for weak coupling, or as a probe of the $\text{MI}_{\pm 1}$ bound state for strong coupling. We argued that these results are relevant for the experimental study of the Bose lattice polaron in an harmonic trap, where the added spatial complexity make those effects possible even for a fixed number of bosons and weak or strong impurity-bath interaction. This work opens several research directions, including the study of bipolaron formation and many-polaron problems in the strong interacting regime.
@@ Acknowledgements.
11footnotetext: See Supplemental Material at [URL will be inserted by publisher] for technical details on the multi-species worm algorithm used for simulation. It includes the implementation of a harmonic constraint to stabilize single-impurity simulations in the grand canonical ensemble. Additional results on the bath’s compressibility, energy gap, and correlation functions are provided. Finite-size scaling analysis of winding numbers and local density shifts across the MI-SF transition are also discussed. The Supplemental Material includes Ref.  [23].